Write an equation in slope-intercept form of the line passing through the given points.
(4, -2) (5, 1)

Answer by mananth(16949)

x1 y1 x2 y2
4 -2 5 1

slope m = (y2-y1)/(x2-x1)
( 1 - -2 )/( 5 - 4 )
( 3 / 1 )
m= 3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 8 , 1 ) in
Y = m x + b
1.00 = 3 * 8 + b
1.00 = 24 + b
b= 1 + -24
b= -23
So the equation will be
Y = 3 x -23
m.ananth@hotmail.ca